Determine whether the given number is a solution to the inequality following it.
0, 4x+9>or=to 5(x-3)

Answer by uma(370)

the given inequality is 4x + 9 > or = 5(x-3)
To check if the given number 0 is a solution, we just plug in x = 0 in the inequality.
==> 4(0) + 9 > or = 5(0-3)
==> 0 + 9 > or = 5(-3)
==> 9 > or = - 15
This statement 9 > -15 is true.
Hence we conclude that 0 is a solution to the given inequality.